OpenStudy (anonymous):

At what depth from the surface of the earth is the value of acceleration due to gravity one-fourth the value at the earth's surface ?

OpenStudy (anonymous):

g'=g[1-(d/R)].. g'=g/4 find d, where d is the depth.

OpenStudy (anonymous):

it'll be 3/4 of r ie 4800 km below the surface.

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

i just gave you the formula.. solve it.

OpenStudy (anonymous):

g' = GM / (R+h)^2

OpenStudy (anonymous):

^^ is this the formula ?

OpenStudy (anonymous):

coz that formula is not available in my book

OpenStudy (anonymous):

nope this is the formula for a height h above the earth. you have to find g at a depth d below the earth.

OpenStudy (anonymous):

g'= g-gd/R

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

this is given in my book for depth

OpenStudy (anonymous):

use this.

OpenStudy (anonymous):

the answer is coming 4785000 is this correct ?

OpenStudy (anonymous):

i havent solved it. its easy calculations. or else use a calculator to check your answer.

OpenStudy (anonymous):

plz tell me is this correct using your on screen calculator plzz

OpenStudy (anonymous):

3/4*R =3/4*6400=3*1600=4800kms.

OpenStudy (anonymous):

how did you get 4785000?

OpenStudy (anonymous):

g' = 9.8/4

OpenStudy (anonymous):

g' = 2.45

OpenStudy (anonymous):

2.45 = 9.8 - 9.8d/6.38x10^6

OpenStudy (anonymous):

its simple put g'=g/4 and g cancels out on both sides.we are left with 1/4=1-d/r.then solve it

OpenStudy (anonymous):

how come g' cancels with g ??

OpenStudy (anonymous):

oh dear put g'=g/4 so g cancels out not g'

OpenStudy (anonymous):

can u draw and show me :)

OpenStudy (anonymous):